# Benchmarking Algorithms for Federated Domain Generalization

### v1

@@ 3.2 Domain-based client heterogeneity by splitting DG datasets
Figure 1: Distinct color refers to distinct domain data, and $\lambda$ is the domain balancing parameter. (a): Train-test domain heterogeneity. (b): Domain split when $C\leq D$ and domain split when $C>D.$ (c): domain split when domain number exceeds client number, i.e., $C\leq D;$ homogeneous ($\lambda=1$), heterogeneous ($\lambda=0.1$), and separation ($\lambda=0$). See Algorithm 1 for domain split when $C>D.$
Algorithm 1 DomainSplit Algorithm
@@
        // Find on average largest domain to split
@@
  // Split samples across the clients
@@
In Algorithm 1, we provide a concrete algorithm for splitting samples from multiple domains across an arbitrary number of clients $C$ while controlling the amount of domain-based client heterogeneity via $\lambda$, ranging from homogeneous clients to domain separation. Our algorithm has two main steps. In Step 1, we assign “primary” domain indices $\mathcal{D}_{c}\subseteq[D]$ to each client $c\in[C]$ depending on $C$ and $D$. If $C\leq D$, the domains are sorted in descending order according to number of sample size and are iteratively assigned to the client $c^{*}$ which currently has the smallest number of training samples (denoted by $\sum_{d^{\prime}\in\mathcal{D}_{c^{*}}}n_{d^{\prime}}$). In this case, the algorithm ensures that no client shares domains with the others but otherwise attempts to balance the total number of training samples between clients. If $C>D$, we first assign the domains one by one to the first $D$ clients. Then, starting from client $c=D+1$, we iteratively assign the current on average largest domain $d^{*}$ to $c,$ accounting for the fact that the domain may already shared by different clients—notationally, on average is represented through dividing by $\sum_{c^{\prime}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]$. In this case, some clients may share one domain, but no client holds two domains simultaneously while otherwise attempting to balance the number of samples across clients as much as possible.
In Step 2, we define the sample counts for each client and domain, denoted $n_{c,d}(\lambda)$ based on the balancing parameter $\lambda\in[0,1]:$ $n_{d,c}(\lambda)=\lambda\frac{n_{d}}{C}+(1\!\!-\!\!\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c}]}n_{d}\,,$ where rounding to integers is carefully handled when not perfectly divisible and where $\mathbbm{1}[\cdot]$ is the indicator function. This is simply a convex combination between homogeneous clients ($\lambda=1$) and domain separation ($\lambda=0$). Given the number of samples per client per domain, we simply sample without replacement from the corresponding domain datasets and build up the client datasets.
3.3 Dataset Selection and Dataset Difficulty Metrics
To ensure a comprehensive evaluation of current methods across different difficulty levels, we have curated a range of datasets with varying complexities. We define two dataset metrics to measure the dataset difficulty with respect to the DG task and with respect to the FL context. For DG difficulty, we compute $R_{\text{DG}},$ the ratio of the ERM performance with and without samples from the test domain (i.e., the former is able to “cheat” by seeing part of test domain samples during training). For FL difficulty, we attempt to isolate the FL effect by computing $R_{\text{FL}}(\lambda),$ the ratio of ERM-based FedAvg $\lambda$ client heterogeneity over centralized ERM on in-domain test samples. These dataset difficulty metrics can be formalized as follows:
$R_{\text{DG}}$ $\triangleq\frac{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}^{\prime}_{\text{DG-test}})}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}}\cup{\mathcal{S}^{\prime\prime}_{\text{DG-test}}},\mathcal{S}^{\prime}_{\text{DG-test}})}$
$R_{\text{FL}}(\lambda)$ $\triangleq\frac{\texttt{FedAvg\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}};\lambda)}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}})},$
where ERM_Perf is the performance of ERM using the first argument as training and the second for test, FedAvg_Perf is similar but with the client heterogeneity parameter $\lambda$, $\mathcal{S}_{\text{DG-train}}$ denotes samples from the train domains $\mathcal{D}_{\text{train}},$ $\mathcal{S}_{\text{DG-test}}$ denotes samples from the test domains $\mathcal{D}_{\text{test}},$ and $\mathcal{S}^{\prime}_{\text{DG-test}}$ and $\mathcal{S}^{\prime\prime}_{\text{DG-test}}$ are $20\%,80\%$ split respectively of $\mathcal{S}_{\text{DG-test}}$. For $R_{\text{FL}}(\lambda)$, we use $\mathcal{S}_{\text{IN-test}}$ (test samples from the train domains) instead of $\mathcal{S}_{\text{DG-test}}$ to isolate the FL effect from DG effect. We choose five datasets in our benchmark: FEMNIST from Caldas et al. (2018), PACS from Gulrajani and Lopez-Paz (2020), and IWildCam, CivilComments and Py150 from Koh et al. (2021). We summarize the statistics and difficulty metrics in Table 2 and provide the rationale of selecting these datasets in the Appendix.
@@ 3.3 Dataset Selection and Dataset Difficulty Metrics
4 Benchmark Experimental Results
In this section, we report the performance of $13$ representative methods from three lines of research on $5$ different datasets, where the FEMINIST results are provided in the appendix given the DG task simplicity (i.e., $R_{\text{DG}}\approx 1$). For each dataset, we fix the total computation and communication rounds for different methods for a fair comparison. We then select the model based on the held-out-domain validation set. After training, we choose the model according to the early-stopping at the communication round which achieves the best held-out-domain performance, and finally we evaluate the performance on the test-domain in Table 3 and Table 4. See Appendix for detailed hyperparameters choices. We also include the results according to in-domain validation early-stopping in the Appendix but the trends are similar. We make the following remarks on the main results from Table 3 and Table 4.
Remark 4.1.
FedAvg with an ERM objective is a strong baseline. Simple FedAvg with an ERM objective is a strong baseline that is challenging to beat across datasets (except for CivilComments), similar to the centralized case stated in DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021). We recommend always including FedAvg as a baseline in all future evaluations.

### v2

@@ A.1 Heterogeneous Partitioning Algorithm and its guarantee
Figure 1: Distinct color refers to distinct domain data, and $\lambda$ is the domain balancing parameter. (a): train-test domain heterogeneity. (b): domain partitioning when $C\leq D$ and domain partitioning when $C>D.$ (c): domain partitioning illustration when $C\leq D;$ homogeneous ($\lambda=1$), heterogeneous ($\lambda=0.1$), and extreme heterogeneous ($\lambda=0$).
We provide the pseudo code for Heterogeneous Partitioning in Alg. 1 and an illustration in Fig. 1. In Fig. 1, (a) represents a multi-domain dataset with $4$ training domains. Each color represents a domain and the height of the bar represents the number of samples in this domain. (b) shows the complete heterogeneity with two scenarios $C<D$ and $C>D$. (c) is a detailed partition result with different interpolating with the $C<D$ case. It can be seen that in the complete heterogeneity case, the partition satisfies constraint $1,2,$ further, the sample size between clients are roughly even. For other cases, True Partition (Constraint) is satisfied and also the sample size between clients are roughly even.
A.2 Other Partition Methods
Algorithm 1 Domain Partition Algorithm
@@ A.2 Other Partition Methods
        // Find on average largest domain to partition
@@
  // Partition samples across the clients
@@
Shards. By adjusting the constant from $2$ to other value, it’s possible to regulate the heterogeneity to a certain extent. However, if the dataset contains imbalance number of samples per domain, shards method could not achieve the complete heterogeneity. For instance, if a dataset contains $2$ domains with $20,30$ data samples respectively. Shard partitioning will not be able to assign each client with data from only one domain.
Dirichlet Partitioning. It draws $q_{c}\sim\text{Dir}(\alpha p),$ where $p$ characterizes a prior domain distribution over $D$ domains for each client $c\in[C].$ $\alpha>0$ is a concentration parameter controlling the heterogeneity among clients. As $\alpha\rightarrow 0$, each client predominantly holds samples of a single label. Conversely, as $\alpha\rightarrow\infty$, the distribution of labels for each client becomes identical. This method is not a desirable partitioning in the following sense. 1) If sampling with replacement, then the datasets across clients may not be disjoint, and some data samples may not be allocated to any client, therefore breaking properties $3$ and $4$ 2) If sampling without replacement, the method is inapplicable when the available data from one domain doesn’t meet the collective demand for this domain’s data from all clients, breaking property $2$
Semantic Partitioning. This method aims to partition dataset based on the feature similarity without domain labels. It first processes the data using a powerful pretrained model. The outputs before the fully connected layer are later used as features. Then, thses features are fit a Gaussian Mixture Model to form $K$ for each clusters for each class $Y$. These $KY$ clusters are then merged iteratively using an optimal bipartite for random label pairs across different class. This partition method is more heuristic and highly rely on the quality of the pretrained model. Furthermore, since there is no hard code label, and the partition is based on KL divergence, this method could not give a complete heterogeneity which violate property $1$. Despite that this method could not guarantee property $1$, it is a good partition method to create “domain"-heterogeneity without having domain labels.
Example 1.
We provide a toy example of partition a dataset containing $5$ training domains with $10,20,30,40,100$ data samples respectively. We illustrate how shards and Dirichlet methods work and highlight that disjointedness flexibility and controllability are violated when our goal is to partition the data into $2$ clients. Shards partitioning might allocation $[10,20,20,0,50]$ samples to the first client, and $[0,0,10,40,50]$ samples to the second client. This lead to the share of domain $3$ and $5$ between the two clients showing the violation of Complete Heterogeneity (Constraint). The Dirichlet partitioning will sample a distribution from a Dirichlet distribution with $\alpha\approx 0$. However, no matter what distribution are sampled, at least $3$ domains will be neglected, leading to the violation of True Partition (Constraint). In contrast, our Heterogeneous Partitioning will allocate $[10,20,30,40,0]$ to the first client and $[0,0,0,0,100]$ to the second client, effectively maintaining Complete Heterogeneity (Constraint), True Partition (Constraint), in addition, the sample size difference is 00 in this case.
Appendix B Current Methods in solving DG
Centralized DG methods
Most work solving in DG lies in the centralize regime. A predominant and effective centralized DG approach is through representation learning. Arjovsky et al. [2019] tries to learn domain-invariant feature by aligning the conditional distribution of $p(Y|X)$ among different domains. Sun and Saenko [2016], and Li et al. [2018] tries to explicitly align the first order and second order momentum in the feature space. There are also methods trying to promote the out-of-distribution generalization by posting constraint on the gradient information among different domains, where Shi et al. [2021] tries to align the gradient among different domains, and Rame et al. [2022] enforces domain invariance in the space of the gradients of the loss. Other approaches include distributionally robust optimization [Sagawa et al., 2019], where this method learns the worst-case distribution scenario of training domains. Furthermore, Zhang et al. [2017] is not specifically design for solving domain generalization, but found to be effective as well when we extrapolate the data from different domains in the training dataset. Xu et al. [2020] introduces domain-mixup which extrapolates examples and features across training domains.
Federated DG methods
Limited research has focused explicitly on solving the Federated DG by design. FedDG Liu et al. [2021] introduced a specific FL paradigm for medical image classification, which involves sharing the amplitude spectrum of images among local clients, violating the privacy protocol. Another approach, FedADG [Zhang et al., 2021], utilizes a generative adversarial network (GAN) framework in FL, where each client contains four models: a featurizer, a classifier, a generator, and a discriminator. FedADG first trains the featurizer and classifier using empirical loss and then trains the generator and discriminator using the GAN approach. However, this method requires training four models simultaneously and tuning numerous hyperparameters, making convergence challenging. A novel aggregation method called Federated Gradient Masking Averaging (FedGMA) [Tenison et al., 2022b] aims to enhance generalization across clients and the global model. FedGMA prioritizes gradient components aligned with the dominant direction across clients while assigning less importance to inconsistent components. FedSR [Nguyen et al., 2022] proposes a simple algorithm that uses two locally-computable regularizers for domain generalization. Given the limited literature on solving domain generalization (DG) in the federated learning (FL) setting, we selected all the aforementioned algorithms.
Algorithm 2 Modified AFL
Initialize $\theta$, $\beta$
  Transmit $\theta$ to all clients $c\in[C].$
  for $t=1,2,\dots,T$ do
     // ——–Update $\theta$——–
     Transmit $\beta$ to all clients $c.$
     for $c=1,2,\dots,C$ do
        Update $\theta_{c}$ using SGD with multiple iteration.
        Submit $\theta_{c}$ to the server.
     end for
     $\theta\leftarrow\sum_{c}\frac{1}{n_{c}}\theta_{c}.$  // Weighted average.
     // ——–Update $\beta$——–
     Transmit $\theta$ to all clients $c$
     for $c=1,2,\dots,C$ do
        Calculate $\ell_{d,c}=\nabla_{\beta_{d}}\mathcal{L}_{c}$
        Submit updated $\ell_{d,c}$.
     end for
     for $d=1,2,\dots,D$ do
        $\beta_{d}\leftarrow\text{Proj}(\beta_{d}+\gamma_{\beta}\sum_{c}\ell_{d,c}).$
     end for
  end for
Output $\theta$
FL methods tackling client heterogeneity
Another line of research in FL aims to guarantee convergence even under client heterogeneity, but these FL-based methods still assume the train and test datasets do not shift (i.e., they do not explicitly tackle train-test heterogeneity of the domain generalization task). We include this line of work because methods considering converging over a heterogeneous dataset might bring better DG ability implicitly. The empirical observation of the statistical challenge in federated learning when local data is non-IID was first made by Zhao et al. [2018]. Several subsequent works have analyzed client heterogeneity by assuming bounded gradients [Yu et al., 2019, Basu et al., 2019, Wang et al., 2019, Li et al., 2019] or bounded gradient dissimilarity [Li et al., 2020], and additionally assuming bounded Hessian dissimilarity [Karimireddy et al., 2020, Khaled et al., 2020, Liang et al., 2019]. From this category, we selected FedProx [Li et al., 2020], which addresses statistical heterogeneity by adding a proximal term to the local subproblem, constraining local updates to be closer to the initial global model. Scaffold [Karimireddy et al., 2020] utilizes variance reduction to account for client heterogeneity.
FL methods tackling fairness
Agnostic Federated Learning (AFL) [Mohri et al., 2019, Du et al., 2021] shares similarities with Domain Generalization in a Federated context. This is evident as both approaches address scenarios where the test distribution diverges from the training distribution. AFL constructs a framework that naturally yields a notion of fairness, where the centralized model is optimized for any target distribution formed by a mixture of the client distributions. Thus, AFL is a good method to evaluate especially when tackling subpopulation shift tasks. AFL introduces a minimax objective which is identical to the GroupDRO [Sagawa et al., 2019],
$\min_{\theta}\max_{\beta\in\Delta_{D}}\mathcal{L}(\theta,\beta)=\sum_{d=1}^{D}\beta_{d}\ell_{d}(\theta),$ (14)
where $\theta$ denotes the model parameter and $\beta$ is a weight vector taking values in the simplex $\Delta_{D}$ of dimension $D.$ AFL introduces an algorithm applying projected gradient ascent on $\beta$ and gradient descent on $\theta$. It is designed for $C=D,$ where they assumes each client as a domain. Further, it requires communication per iteration. We made the following two modifications to accommodate the general $C\neq D$ cases and expensive communication cost:
  1. 1.
    To allow $C\neq D,$ we construct new objective as the following:
    $\min_{\theta}\max_{\beta\in\Delta_{D}}\mathcal{L}(\theta,\beta)=\sum_{c=1}^{C}\mathcal{L}_{c}(\theta,\beta)=\sum_{c=1}^{C}\sum_{d=1}^{D}\beta_{d}\ell_{d,c}(\theta).$ (15)
  2. 2.
    To reduce communication cost, we allow $\theta$ to be updated multiple iterations locally per communication. Further, we maintain the update of $\beta$ on the central server, as in AFL, given that it requires projection of the global updates onto the simplex. This projection is not equivalent to averaging the locally projected updates of $\beta.$
Modified AFL avoids the frequent communication compared to AFL.
Appendix C Datasets and Difficulty Metric
C.1 Dataset Introduction
In this section, we introduce the datasets we used in our experiments, and the partition method we used to build heterogeneous datasets in the training and testing phase as well as the heterogeneous local training datasets among clients in the FL.
FEMNIST
PACS
IWildCam
It is a real-world image classification dataset based on wild animal camera traps around the world, where each camera represents a domain. It contains 243 training domains, 32 validation and 48 test domains. Usually people cares about rare speicies, thus we utilize the macro F1 score as the evaluation metric instead of standard accuracy, as recommended in the original dataset’s reference [Koh et al., 2021]. The $R_{\text{DG}}=0.449$ makes it a very challenging dataset for domain generalization.
CelebA
CelebA (Celebrity Attribute) [Liu et al., 2015] is a large-scale face attributes dataset. It’s one of the most popular datasets used in the facial recognition and facial attribute recognition domains. We use a variant of the original CelebA dataset from [Sagawa et al., 2019] using hair color as the classification target and gender as domain labels. This forms a subpopulation shift task which is a special kind of domain generalization.
Camelyon17
It is a medical dataset consisting of a set of whole-slide images from multiple institutions (as domain label) to detect breast cancer. It consist $3$ training domains, $1$ validation domain and $1$ test domain. We partition the data following wilds benchmark [Koh et al., 2021].
CivilComments
It is a real-world binary classification text-based dataset formed from the comments from different demographic groups of people, containing $8$ demographic group. The goal is to judge whether the text is malicious or not, which is also a subpopulation shift dataset.
Py150
@@ C.1 Dataset Introduction
C.2 Dataset partition Setup
For each dataset, we first partition the dataset into $5$ categories, namely training dataset, in-domain validation dataset, in-domain test dataset, held-out validation dataset and test domain dataset. For FEMNIST and PACS dataset, we use Cartoon and Sketch as training domain, Art-painting as held-out domain, Painting as test domain. For training domain, we partition $10\%,10\%$ of the total training domain datasets as in-domain validation dataset and in-domain test dataset respectively. For IWildCam, CivilComments and Py150, we directly apply the Wilds official partitions.
Appendix D Benchmark Experimental Setting
D.1 Model Structure
In this benchmark, for image-based datasets, we use ResNet-50 He et al. [2016] dataset. For CivilComments and Py150 dataset, we choose DistilBERT Sanh et al. [2020] and CodeGPT Radford et al. [2019] respectively as recommended by Wilds [Koh et al., 2021].
D.2 Model Selection
D.3 Early Stopping
We conduct early stopping using the held-out validation dataset in our evaluation. For each dataset and method, We first run certain communication rounds, then we select the model parameters which achieves the best performance on the validation set. We report the held-out validation dataset in the main paper, and we report the results using the in-domain validation set in Appendix F.
D.4 Hyperparameters
In this section, we present the hyperparameters selected for the evaluation. We grid search 8 times of run starting with learning rate same as ERM and other hyperparameters from the methods’ original parameters. We than select the hyperparameter based on the best performance on the held-out domain validation set. Please refer to the Table 7 to review all hyperparameters.
Table 7: Hyperparameters used in the evaluation.
Dataset Name PACS CelebA Camelyon17 FEMNIST IWildCam CivilComments Py150
Total Communication $80$ $20$ $20$ $40$ $50$ $10$ $10$
ERM lr $3\times 10^{-5}$ $3\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$
IRM lr $3\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-5}$ $1\times 10^{-6}$ $1\times 10^{-6}$
penalty weight $1000$ $1000$ $1000$ $100$ $100$ $10$ $1000$
Fish lr $3\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-6}$ $1\times 10^{-6}$
meta-lr $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $1$
Mixup lr $3\times 10^{-5}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-5}$
$\alpha$ $0.2$ $0.2$ $0.1$ $0.2$ $0.2$
MMD lr $3\times 10^{-5}$ $0.0001$ $1\times 10^{-3}$ $1\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-6}$ $1\times 10^{-5}$
penalty weight $1$ $1$ $0.1$ $1$ $1$ $10$ $1$
Coral lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-6}$ $5\times 10^{-5}$
penalty weight $1$ $1$ $10$ $10$ $1$ $1$ $1$
GroupDRO lr $3\times 10^{-5}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
group lr $0.01$ $0.05$ $0.01$ $1\times 10^{-5}$ $1\times 10^{-5}$ $0.05$ $0.005$
FedProx lr $3\times 10^{-5}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-6}$
$\mu$ $1\times 10^{-3}$ $0.1$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.1$ $0.01$ $1\times 10^{-3}$
Scaffold lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $3\times 10^{-6}$ $8\times 10^{-6}$
AFL lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $3\times 10^{-6}$ $8\times 10^{-6}$
FedDG lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
$\lambda$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$
FedADG Classifier lr $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Gen lr $7\times 10^{-4}$ $5\times 10^{-4}$ $7\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$
Disc lr $7\times 10^{-4}$ $5\times 10^{-4}$ $7\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$
$\alpha$ $0.15$ $0.25$ $0.1$ $0.05$ $0.2$
FedSR lr $1\times 10^{-5}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-4}$ $3\times 10^{-5}$ $5\times 10^{-6}$ $5\times 10^{-6}$
l2 regularizer $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
cmi regularizer $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
FedGMA lr $3\times 10^{-5}$ $1\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-6}$ $1\times 10^{-5}$
mask-threshold $0.1$ $0.1$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$
Appendix E Additional FL-specific challenges for domain generalization
As mentioned in subsection 4.3, we also include some deeper exploration over the effect of number of clients and communication frequency, which are unique to the FL regime.
i) Massive number of clients: In this experiment, we explore the performance of different algorithms when the number of clients $C$ increases on PACS. We fix the communication rounds $50$ and the local number of epoch is $1$ (synchronizing the models every epoch). Fig. 2 plots the held-out DG test accuracy versus number of clients for different levels of data heterogeneity. The following comments are in order: given communication budget, 1) current domain generalization methods all degrade a lot after $C\geq 10$, while the performance ERM and FedDG maintain relatively unchanged as the clients number increases given communication budget. FedADG and FedSR are are sensitive to the clients number, and they both fail after $C\geq 20.$ 2) Even in the simplest homogeneous setting $\lambda=1,$ where each local client has i.i.d training data, current domain generalization methods IRM, FISH, Mixup, MMD, Coral, GroupDRO work poorly in the existence of large clients number, this means new methods are needed for DG in FL context when data are stored among massive number of clients.
Figure 2: Performance based on accuracy versus the number of clients across the PACS, CelebA, and Camelyon17 datasets.
Appendix F Supplementary Results
In the main paper, we provide experiments results using held-out validation early stopping. Here we report the results using in-domain validation set for reference. We also report the convergence curve for each methods on each dataset for reference. We observe that under most of cases, the held-out validation gives us a better model. Thus, we recommend using held-out validation set to perform early stopping when we can access multiple training domains. However, if the training dataset contains only 2-3 domains, we should consider using in-domain validation set.
More results on PACS and IWildCam dataset.
Results on FEMNIST dataset.
As mentioned in the main paper, we include the FEMNIST dataset here for reference. It is evident from Table 11 that $\lambda$ does not significantly influence the final domain generalization (DG) accuracy, and whether using in-domain validation or held-out-domain validation does not impact the final DG accuracy as well. This suggests a lack of statistical heterogeneity across different domains. Furthermore, we observe that changing $\lambda$ does not significantly affect the convergence. Most of the results do not converge to the performance level of the centralized counterpart. This discrepancy arises from the challenge posed by the large number of clients, where $R_{\text{FL}}=0.980<1$.
Figure 7: Accuracy versus communication rounds for Camelyon17; total clients and training domains $(K,M)=(100,4);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
Figure 8: Accuracy versus communication rounds for CelebA; total clients and training domains $(K,M)=(100,2);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$